OpenStudy (anonymous):

Simplify. 4-8*3/4+1/4

OpenStudy (camerondoherty):

\[-\frac{ 7 }{ 4 }\]Would you like me to show you how i did it?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

First division, then multiply, then add or subtraction as you wish..

OpenStudy (camerondoherty):

Step 1 : Simplify \[-2+\frac{ 1 }{ 4 }\] Add a fraction to a whole Rewrite the whole as a fraction using 4 as the denominator : \[2=\frac{ -2 }{ 1 }=\frac{ -2\times4 }{ 4 }\] Equivalent fraction : The fraction looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Add up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: \[\frac{ -2\times4+1 }{ 4 }=\frac{ -7 }{ 4 }\] Final result : \[\frac{ -7 }{ 4 }\]